### v1

# TWIST: Training-free and Label-free Short Text Clustering through Iterative Vector Updating with LLMs

Abstract
In this paper, we propose a training-free and label-free method for short text clustering that can be used on top of any existing embedder. In the context of customer-facing chatbots, companies are dealing with large amounts of user utterances that need to be clustered according to their intent. In these commercial settings, no labeled data is typically available, and the number of clusters is not known. Our method is based on iterative vector updating: it constructs sparse vectors based on representative texts, and then iteratively refines them through LLM guidance. Our method achieves comparable or superior results to state-of-the-art methods that use contrastive learning, but without assuming prior knowledge of clusters or labels. Experiments on diverse datasets and smaller LLMs show that our method is model agnostic and can be applied to any embedder, with relatively small LLMs, and different clustering methods. We also show that our method scales to large datasets, reducing the computational cost of the LLM. These low-resource, adaptable settings and the scalability of our method make it more aligned with real-world scenarios than existing clustering methods. Our source code is available here: https://anonymous.4open.science/r/sparse_vector-F4C9/README.md
1 Introduction
Short text clustering is an NLP task that automatically groups unlabeled data into clusters and is widely used in practical applications for text classification and new category detection. For example, in conversational understanding, user intents in real-world scenarios are dynamic and continuously evolving (Liang et al., 2024c), and relying on expert labeling can be costly. The short-text clustering task enables companies to efficiently organize tens of thousands of client requests in real time. Other downstream tasks, such as mining topics and opinions from online platforms, also rely on effective clustering (Wu et al., 2024).
The clustering task is often approached as a Generalized Category Discovery (GCD) scenario (Vaze et al., 2022), where limited labeled data and substantial unlabeled data are used to cluster all unlabeled samples, which include both known and unknown classes. Recently, other studies have explored an emerging data scenario, where no labels are provided at all (De Raedt et al., 2023; Viswanathan et al., 2024b; Zhang et al., 2023; Lin et al., 2025). This scenario reduces reliance on human annotation and shows greater flexibility.
A common approach for the emerging data scenarios is to train an embedder with contrastive learning to improve text representations (Zhang et al., 2021). With the advancement of LLMs, training-free approaches have also been proposed (De Raedt et al., 2023). The emerging data scenario has attracted increasing attention in recent research (Zhang et al., 2023; Feng et al., 2024; Lin et al., 2025). Despite the progress of prior work, these label-free approaches still face certain limitations and leave room for further improvement. First, methods involving LLM feedback often assume known number of clusters during method development (Zhang et al., 2023; Feng et al., 2024), which is difficult to estimate in real-world settings. Second, training-free methods have mainly been explored with large, proprietary models (Viswanathan et al., 2024b; De Raedt et al., 2023), which can be costly. Third, the scalability of these methods remains a concern. Motivated by these challenges, we propose a training-free approach that requires only a smaller LLM, alleviates the known-cluster assumption, and is scalable. The core idea of our method is to construct new representations (referred to as sparse vectors in this paper) such that texts belonging to the same cluster become more similar. Our method consists of two stages: we first construct sparse representations based on representative texts, and then iteratively refine them by combining LLM guidance with pre-trained embeddings.
The contributions of our method are as follows:
  • Inclusive and easy to optimize: Builds on existing embedders and requires only some hyperparameter tuning.
  • Flexible: Does not assume a known number of clusters, nor require labeled data, contrastive learning on the task specific datasets, or fine-tuning, enabling easier adaptation to new domains.
  • Low-resource: Even with small LLMs and embedders, the results are already competitive with methods that rely on labeled data or fine-tuning.
  • Scalable: Does not require LLM feedback for every text, unlike many previous studies.
2 Related Work
Short text clustering
Generalized Category Discovery (GCD) is often studied for this task (Liang et al., 2024b; Liang and Liao, 2023; Zou et al., 2025b). These works use limited labeled text and substantial unlabeled text to learn clustering representations. However, these studies use train–test splits within a single dataset, which implicitly assumes that new categories are similar to the known ones. In practice, emerging data may show topic drift. Recently, with the advancement of LLMs, many studies have explored the emerging data scenario (Zhang et al., 2023; Viswanathan et al., 2024b; Lin et al., 2025), where no labels are available. We follow this unsupervised scenario.
Boosting Unsupervised Short Text Clustering with LLM-Guided Contrastive Learning
Contrastive learning is commonly used in unsupervised short text clustering (Zhang et al., 2021, 2023; Liang et al., 2024a). Zhang et al. (2023) use LLMs to construct hard triplets to derive training pairs, which are then used to fine-tune a small embedder. Liang et al. (2024a) use LLMs to refine texts by generating descriptors, then using small language models for contrastive learning to align the embeddings of texts and their descriptors. However, these methods either rely on knowing the number of clusters, proprietary models, or fine-tuning. Our method does not have these requirements.
Training-Free Methods for Unsupervised Short Text Clustering
Training-free approaches use a frozen pre-trained embedder: De Raedt et al. (2023) select prototypical utterances and then use an LLM to generate a set of descriptive utterance labels for these prototypes. The labels and texts are then encoded to form better representations. Viswanathan et al. (2024a) use LLMs to improve texts by generating key phrases for each text, which are then separately encoded into embeddings and concatenated for clustering. Lin et al. (2025) use LLMs for selection rather than text enrichment. They first encode all of the texts, and then use the LLM to select similar pairs for each text. Each text embedding is then obtained by average pooling the embeddings of the text and its similar pairs. These methods have in common that they refine original embeddings to improve clustering. In contrast, we construct a new vector using LLM feedback. Feng et al. (2024) use LLMs to refine edge points. However, the assumption of a known number of clusters can still be a core problem in real-world applications.
3 Theoretic formulation and assumptions
3.1 Problem definition
In our work, we consider two variations of the short text clustering task: (i) the emerging data scenario and (ii) the GCD scenario. In the emerging data scenario (Lin et al., 2025), the only available resource is a collection of unlabeled texts $D^{test}\triangleq\{x_{i}\}_{i=1}^{N}$, where each $x_{i}\in D^{test}$ corresponds to a short text and $N$ is the total number of samples. The dataset is used as a whole for clustering and evaluation, without any train-test split. In contrast, the GCD scenario has access to additional data, typically including some labeled samples (Zou et al., 2025b). In this work, we adapt this setting by considering only an additional unlabeled dataset $D^{u}\triangleq\{x_{i}\}_{i=1}^{N_{u}}$, in line with the goal of low-resource. The GCD scenario is included for the purpose of testing whether extra data improves clustering and generalizability of our method.
For both scenarios, the task is to partition $D^{test}$ into clusters $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$, where $\hat{K}$ is the number of predicted clusters. Note that $\sum_{l=1}^{\hat{K}}|\hat{S}_{l}|=N$. The goal is to ensure that the clustering results $\{\hat{S}_{l}\}_{l=1}^{\hat{K}}$ are as close as possible to the true partition $\{S_{l}\}_{l=1}^{K}$, where $K$ is the number of unique clusters in the ground truth. Importantly, unlike many existing studies (De Raedt et al., 2023; Feng et al., 2024; Zou et al., 2025b), which assume that $K$ is known in advance, we make no such assumption.
For notation convenience, we define the full dataset as $D\triangleq D^{test}\cup D^{u}$ to simplify the explanation of our proposed method. Note that $D=D^{test}$ in the emerging data scenario.
3.2 Constructing the Sparse Vector
Basic Assumption
Considering the fact that texts belonging to the same cluster are similar to each other, we represent each text as a sparse bag-of-texts vector that directly encodes its similarity to other texts. Formally, for any $x_{i},x_{j}\in D$ with $i\neq j$:
  • $x_{i}$ and $x_{j}$ are similar if $x_{i},x_{j}\in S_{k}$ for some cluster $S_{k}$.
  • $x_{i}$ and $x_{j}$ are not similar if $x_{i}\in S_{k}$ and $x_{j}\in S_{l}$ with $k\neq l$.
Construction of Bag-of-Texts
Since similarity is determined by clusters, it is not necessary to use all texts in $D$ to represent sparse vectors; instead, a representative subset of texts can constitute our bag of texts, denoted as $D^{r}\subseteq D$, where $|D^{r}|\triangleq d$. The set of representative texts $D^{r}$ should contain at least one element from each cluster. That is, for any cluster $S\in\{S_{l}\}^{K}_{l=1}$, there exists at least one text $x\in D^{r}$ such that $x\in S$.
Therefore, to cover all possible clusters, the minimum size of $d$ is K1K-1. Since $K$ is unknown, the most practical approach is to choose a relatively large $d$. For each $x\in D$, we define a $d$-diemnsional vector $\mathbf{z}$ assigned to the representative texts, where the $j$-th coordinate takes the value 1 if $x$ is similar to the $j$-th representative text, and 0 otherwise.
Proposition 1.
Given the assumption, for any $x_{i},x_{j}\in D$ with $i\neq j$ that belong to the same cluster $S_{k}$, the corresponding sparse vectors $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ will be identical.
The proposition is straightforward, so we omit the proof.
Adaptation to real-world practice
Although the proposition holds in the idealized setting, in practice similarity is typically determined using LLMs, which may introduce errors, and scanning the entire dataset for each $x$ is computationally expensive. To address these, we construct a sparse vector $\mathbf{z}$ in a way that is adaptable to real-world practice: for each $x\in D$, instead of assigning fixed value in $\{0,1\}$ based on representative texts, we propose an updating procedure that accounts for uncertainty. In addition, to address limited computational resources, we use an existing embedder to identify a relatively small set of $m$ similar candidates that are likely to belong to the same category, as empirically shown in previous research (Lin et al., 2025; Zou et al., 2025a). The details of the method will be discussed in the next section.
4 Computational Method
Our method aims to construct sparse vectors that capture the similarity of texts within the same cluster. It consists of two stages. In the initial stage, we select representative texts, which are then used to construct a sparse vector for each text. In the second stage, we iteratively update each text’s sparse vector by considering its similar texts until convergence, illustrated in Figure 1.
Unlike many prior studies that assume the number of clusters $K$ is known, we consider a more practical setting in which $K$ is unknown, and thus approximate the true number of clusters by the number of representative texts, $d$. This is less challenging than estimating $K$, as it only requires ensuring that $d$ is large enough to cover $K$ clusters.
Figure 1: Construction of Sparse vectors. Initial stage: We first encode each text in $D$ using a pretrained embedder and partition them into $d$ clusters. We then select the medoid of each cluster as the representative text and construct one-hot vectors for these representatives. The remaining text vectors are obtained based on LLM selection. Iterative stage: We update all of the vectors iteratively until convergence.
4.1 Initial stage: construction of sparse representation
To build sparse vector representations, we first need to create a set of representative samples $D^{r}$. To this end, we first encode each text using an existing embedder $F$ to obtain pre-trained embeddings:
$\mathbf{x}_{i}=F(x_{i}).$
Next, we apply a clustering algorithm to partition these embeddings into $d$ clusters.
We use agglomerative clustering (Murtagh and Legendre, 2014) as it produces deterministic results, which facilitates reproducibility. For each initial cluster , we derive the medoid. We refer to the set of medoids as the representative text, denoted as:
$D^{r}\triangleq\{x_{c_{1}},\dots,x_{c_{d}}\}.$
and let $\mathbf{X}^{r}$ be $D^{r}$ corresponding pre-trained embeddings, where $\mathbf{X}^{r}=\{\mathbf{x}_{i}:x_{i}\in D^{r}\}$. We then transform the representative texts into one-hot vectors, which serve as the basis vectors for the sparse representation.
For the remaining texts $x_{j}\in D\setminus D^{r}$, we construct their sparse vectors using this basis. For each $x_{j}$, we compute the pairwise cosine distances between its embedding $\mathbf{x}_{j}$ and the medoid embeddings in $\mathbf{X}^{r}$. By sorting these distances, we select the top $m$ closest representative texts to $x_{j}$, denoted as $x_{j1},\dots,x_{jm}$. Then, we use an LLM to identify the similar text(s) to $x_{j}$ among these representative texts, denoted as $M_{j}\triangleq LLM(x,x_{j1},\dots,x_{jm})\in\mathcal{P}(\{x_{j1},\dots,x_{jm}\})\cup\{\text{None}\}$, where $\mathcal{P}(\cdot)$ represents the power set, i.e., all possible subsets of the candidate texts, and ‘None’ indicates that no representative text is selected. If the LLM output is not None, the sparse vector of $x_{j}$ is then derived by averaging the sparse vectors of the texts in $M_{j}$, denoted as $\mathbf{z}_{j}$. Instead of assigning a fixed value of 1 to the selected dimension, we average the vectors to capture how strongly each representative text contributes to the feature. We then apply L2 normalization to ensure a consistent scale for the iterative stage. If the LLM output is None, we make this text into a representative text itself and increase the dimensionality of all sparse vectors by one and construct its corresponding one-hot vectors.
Note: if we set $d=N$, all texts will be representative texts and one-hot encoded.
4.2 Iterative stage: sparse vectors refinements
In this stage, we update the sparse vectors of all texts using the dynamic subset of $m$ candidates.
At each iteration step $t$ (with the initial stage being $t=0$), for each $x_{j}\in D$, we first concatenate its pre-trained embedding and sparse vector: $\mathbf{u}_{j}^{t}\triangleq\mathbf{x}_{j}\oplus\mathbf{z}_{j}^{t}$. We then compute the pairwise cosine distances between $\mathbf{u}_{j}^{t}$ and the concatenated representations of the all texts $\mathbf{U}^{t}$. By sorting these distances, we select the top $m$ closest texts to $x_{j}$, denoted as $x^{t}_{j1},\dots,x^{t}_{jm}$. The concatenation ensures that the pre-trained embedding guides the sparse vectors. Also, the reason we do iterative updating instead of just once is to gain more stable sparse vectors.
Next, we use an LLM to identify the subset of the texts most similar to $x_{j}$, denoted as $M^{t}_{j}\triangleq LLM(x_{j},x^{t}_{j1},\dots,x^{t}_{jm})\in\mathcal{P}(\{x^{t}_{j1},\dots,x^{t}_{jm}\})\cup\{\text{None}\}$. If the LLM output is None, we fix sparse vectors, i.e., $\mathbf{z}_{j}^{t+1}=\mathbf{z}_{j}^{t}$. Otherwise, we update the sparse vectors by averaging the sparse vector of the texts in $M^{t}_{j}$. We then apply L2 normalization, as explained previously.
For $x_{j}$, the update process is ended if the sparse vector at iteration $t+1$ is almost identical to those at iteration $t$, as we define by a cosine similarity greater than 0.99. For efficiency, we set $m=30$ and cap the maximum number of iterations at 10.
5 Experimental Settings
Emerging data $K$ $|D^{test}|$ $|D^{u}|$
Bank77 77 3,080 -
Clinc150 150 4,500 -
Mtop 102 4,386 -
Massive 59 2,974 -
GoEmo 27 5,940 -
GCD $K$ $|D^{test}|$ $|D^{u}|$
Bank77 77 3,080 2,700
Clinc150 150 2,250 5,400
Stackoverflow 20 1,000 5,400
Table 1: Dataset Statistics in two scenarios. Emerging data scenario: we use the same settings as Zhang et al. (2023); Lin et al. (2025). GCD scenario: for fair comparison, we evaluate on the same test sets as De Raedt et al. (2023); Liang et al. (2024a); Zou et al. (2025b), but we only use $30\%$ of the original training set, which contains unlabeled data only. Further details are provided in the Appendix A.
5.1 Datasets and models
We use different topics of datasets for our experiments, including the intent datasets Bank77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), Mtop (Li et al., 2020), Massive (Fitzgerald et al., 2023); the emotion dataset GoEmo (Demszky et al., 2020), and the community QA dataset Stackoverflow (Xu et al., 2015) for our experiments. Statistics are provided in Table 1.
For the emerging data scenario, we use the embedders e5-large (Wang et al., 2022) and instructor-large (Su et al., 2022) to compare with previous studies (Zhang et al., 2023; Feng et al., 2024). We also experiment with bert-base-uncased (Devlin et al., 2018), all-MiniLM-L6-v2 (Reimers and Gurevych, 2019). For the GCD scenario, we use all-MiniLM-L6-v2 (22.7M paramters) as the embedder, since most prior studies (An et al., 2024; Zou et al., 2025b) fine-tune bert-base (110M paramters). Using all-MiniLM-L6-v2 allows us to minimize the impact of model size on our experiments.
For the LLM experiments, the main results are obtained using gemma-2-9b-it (Rivière et al., 2024). Llama3.1-8B-Instruct AI@Meta (2024) and Qwen/Qwen3-8B (Team, 2025) are used for hyperparameter search or further analysis. An A100 GPU or an H100 GPU were used throughout the experiments.
5.2 Prompts used
Prompt for LLM
The prompts we used are shown in Table 2. It consists of three parts: task instruction, answer format, and task.
CLINC150, Bank77, Mtop, and Massive
Task Instructions:
Compare each Candidate Utterance to the Target Utterance. Select only Candidates with the same intent as the Target Utterance. Intent refers to the request or the purpose the user wants to achieve.
GoEmo
Task Instructions:
Compare each Candidate Utterance to the Target Utterance. Select only Candidates with the same emotion as the Target Utterance.
Stackoverflow
Task Instructions:
Compare each Candidate Question to the Target Question. Select only Candidates with the same main programming framework, language, tool, or concept as the Target Question.
Answer Format:
Only provide the final selection of Candidate Utterances by listing their numbers if they match the Target Utterance intent or request. 1. If Candidates 3, 4, and 9 match, write: The Candidate utterance number(s): 3, 4, 9 2. If no Candidates match, write: The Candidate utterance number(s): none Note: Stick to the answer format and avoid providing extra explanations.
Task:
Target Utterance: {sentence 1}
Candidate Utterances:
1. {sentence 1}
m. {sentence m}
Table 2: Task Instructions Prompt. Note: The boldface used here is for readability; it is not used in the prompt. If the dataset is Stackoverflow, we use ‘question’ instead of ‘utterance’ in the Answer format.
5.3 Hyperparameter setting
The main hyperparameter in our approach is the initial dimension (number of representatives) of the sparse vector $d$, which we set to 2,048 based on validation results on two external datasets: ArxivS2S and Reddit (Muennighoff et al., 2023). We use external datasets to prevent overfitting on a specific dataset. See details in Appendix B. Qwen3-8B (as LLM) and All-MiniLM-L6-v2 (as embedder) are used in the hyperparameter search.
5.4 Evaluation
The emerging and GCD scenarios are evaluated under two conditions: with HDBSCAN (McInnes et al., 2017) and K-means (Lloyd, 1982).
To reflect real-world settings with unknown $K$, we apply HDBSCAN (McInnes et al., 2017) to partition $D^{test}$ into $\hat{K}$ clusters. We set min_samples = 1 to reduce noise. The noise points are assigned to the nearest cluster centroid. We perform an incremental grid search over min_cluster_size values of $10,15,\ldots,50$, stopping early if the silhouette score does not improve (Rousseeuw, 1987; Gung et al., 2023).
K-means clustering is the dominant evaluation practice because many studies assume a known cluster number $K$ and incorporate it in their method pipelines Zhang et al. (2023); De Raedt et al. (2023); Liang et al. (2024b); Feng et al. (2024); Zou et al. (2025b). Following their convention, the number of clusters for K-means is set to match the number of ground-truth cluster number $K$.
To reduce cost, we follow (Zhang et al., 2023) and perform a single pipeline run per dataset using a fixed seed to obtain a sparse vector for each text. Most steps are deterministic. The main source of variability comes from splitting candidate examples across multiple LLM calls (e.g., three calls with 10 examples instead of one with 30), done only to save time and GPU memory. LLM outputs are generated via greedy decoding, without sampling. For evaluation, K-means is run over 5 different seeds and averaged. HDBSCAN, which is near-deterministic, is run once. After we derive the clusters, we apply standard clustering metrics for evaluation. These metrics include normalized mutual information (NMI) and clustering accuracy (Acc) (Rand, 1971; Meilă, 2007; Huang et al., 2014; Gung et al., 2023).
5.5 Baselines
Emerging Data Scenario
We compare against SCCL (Zhang et al., 2021), ClusterLLM, (Zhang et al., 2023), IDAS De Raedt et al. (2023), LLMedgeRefine (Feng et al., 2024), and SPILL (Lin et al., 2025). Table 3 shows that our method makes the fewest assumptions and has the lowest requirements. SPILL is the only method under the same assumptions. However, we still include ClusterLLM, IDAS, and LLMedgeRefine, which uses additional information – the known $K$ – in its development pipeline, to provide a more comprehensive comparison.
GCD Scenario
We compare against IDAS De Raedt et al. (2023), Loop (An et al., 2024), ALUP Liang et al. (2024b), IntentGPT Rodriguez et al. (2024) and Glean Zou et al. (2025b). Note that IntentGPT and IDAS can also be implemented only using unlabeled data. Table 3 shows that our method again makes the fewest assumptions and has the lowest requirements.
Emerging data FT known $K$ Labeled data
TWIST (Ours) -
SCCL (Zhang et al., 2021) $\checkmark$ -
ClusterLLM (Zhang et al., 2023) $\checkmark$ $\checkmark$ -
LLMEdgeRefine (Feng et al., 2024) $\checkmark$ -
IDAS (De Raedt et al., 2023) $\checkmark$ -
SPILL (Lin et al., 2025) -
GCD
TWIST (Ours)
Loop (An et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$
ALUP (Liang et al., 2024b) $\checkmark$ $\checkmark$ $\checkmark$
IntentGPT (Rodriguez et al., 2024)
Glean (Zou et al., 2025b) $\checkmark$ $\checkmark$ $\checkmark$
Table 3: Baselines settings. FT: Finetuning. known $K$: Assuming known $K$ during the method pipeline development. Labeled data: Labeled data is required.
6 Results
6.1 Main results
Emerging data scenario
Table 4 shows the HDBSCAN results. Our approach substantially improves performance compared to directly using the embedder. [^1] The estimated number of clusters, $\hat{K}$, is provided in Appendix D. Moreover, the results from BERT (non-contrastive) also improve with our approach. For the K-means results, Table 5 shows that our approach generally outperforms IDAS and ClusterLLM on both metrics across different embedders, and achieves comparable results to LLMEdgeRefine. This indicates that, even without assuming prior knowledge of $K$ during method development – as the compared methods do – our approach remains effective and better reflects real-world practice. While SPILL achieves slightly higher NMI, our method attains substantially higher ACC.
Lastly, we found that the HDBSCAN results overall are similar to those of K-means, indicating that our method remains robust even without knowing $K$.
Bank 77 Clinc150 Mtop Massive GoEmo Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Bert
Plain 0.84 1.69 73.03 50.87 63.95 40.97 2.86 8.18 1.35 12.05 28.41 22.75
TWIST (Ours) 76.33 60.23 93.23 85.91 73.86 47.47 72.02 62.32 21.01 14.19 67.29 54.02
MiniLM
Plain 78.78 60.39 86.79 67.71 67.30 55.91 62.55 48.15 11.31 18.62 61.35 50.16
TWIST (Ours) 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 70.46 60.57
E5
Plain 78.68 62.18 87.57 65.62 63.45 53.79 1.15 7.74 15.43 17.70 49.26 41.41
TWIST (Ours) 84.86 72.24 93.41 86.89 72.53 44.32 75.85 65.12 27.77 29.03 70.88 59.52
Instructor
Plain 82.50 67.34 92.29 79.27 67.63 54.31 67.99 53.54 0.25 11.48 62.13 53.19
TWIST (Ours) 85.63 73.77 93.64 84.62 74.06 46.81 76.52 64.07 31.64 33.65 72.30 60.58
Table 4: Five-benchmark results (Emerging Data Scenario, HDBSCAN). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder.
Bank 77 Clinc150 Mtop Massive GoEmo Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Bert
Plain 50.72 (0.42) 29.07 (0.85) 72.90 (0.60) 49.48 (0.83) 62.24 (0.46) 26.78 (0.32) 49.12 (0.64) 30.49 (1.00) 10.77 (0.46) 12.90 (0.37) 49.15 29.75
TWIST (Ours) 76.23 (0.14) 60.49 (0.72) 93.25 (0.07) 86.43 (0.65) 71.79 (0.20) 36.41 (0.20) 72.00 (0.13) 60.56 (0.61) 18.66 (0.10) 18.76 (0.29) 66.38 52.53
MiniLM
Plain 79.08 (0.80) 61.27 (1.36) 89.46 (0.19) 73.73 (0.95) 67.41 (0.53) 31.40 (1.62) 70.37 (0.62) 54.04 (1.28) 10.84 (0.16) 13.89 (0.63) 63.43 46.87
TWIST (Ours) 85.57 (0.06) 74.07 (0.71) 93.65 (0.07) 86.35 (0.34) 71.92 (0.22) 37.83 (0.72) 76.18 (0.26) 66.15 (0.58) 20.69 (0.09) 24.44 (0.37) 69.60 57.77
E5
Plain 77.86 (0.35) 60.88 (0.89) 91.09 (0.17) 75.70 (0.36) 71.17 (0.28) 33.04 (0.74) 71.70 (0.84) 53.97 (2.00) 21.13 (0.44) 21.93 (0.60) 66.59 49.10
SCCL 77.34 (0.62) 63.60 (1.37) 91.89 (0.49) 77.96 (1.78) 70.42 ((0.34) 33.82 ((1.07) 71.57 (0.89) 54.48 (1.80) 20.05 (0.63) 22.03 (0.69) 66.25 50.38
ClusterLLM 84.16 (0.36) 70.13 (1.34) 92.92 (0.29) 80.48 (0.93) 74.46 (0.11) 37.22 (1.18) 74.39 (0.21) 56.08 (1.01) 22.23 (0.17) 22.22 (1.15) 69.63 53.23
SPILL 83.56 (0.47) 70.25 (1.56) 92.93 (0.08) 83.18 (0.78) 71.77 (0.35) 36.83 (1.10) 75.40 (0.51) 60.28 (1.81) 25.14 (0.36) 24.82 (0.86) 69.76 55.07
TWIST (Ours) 85.09 (0.19) 74.23 (0.88) 93.31 (0.04) 85.30 (0.58) 71.57 (0.18) 37.32 (1.67) 75.72 (0.22) 62.93 (1.28) 26.58 (0.23) 33.20 (0.54) 70.45 58.60
Instructor
Plain 82.42 (0.68) 65.49 (2.41) 93.27 (0.32) 81.68 (1.31) 71.63 (0.39) 33.14 (0.52) 74.02 (0.54) 54.62 (2.06) 29.18 (0.76) 32.51 (1.51) 70.10 53.49
SCCL 81.77 (1.36) 65.48 (1.36) 92.94 (0.44) 80.85 (0.74) 73.52 (0.38) 34.28 (0.58) 73.90 ((0.36) 54.10 ((1.05) 30.54 (0.68) 34.33 (0.86) 70.53 53.81
ClusterLLM 85.15 (0.41) 71.20 (1.59) 94.00 (0.21) 83.80 (0.41) 73.83 (0.79) 35.04 (0.97) 77.64 (0.21) 60.69 (0.96) 23.89 (0.68) 26.75 (1.76) 70.90 55.50
LLMEdgeRefine - (0.00) - (0.00) 94.86 (0.00) 86.77 (0.00) 72.92 (0.00) 46.00 (0.00) 76.66 (0.00) 63.42 (0.00) 29.74 (0.00) 34.76 (0.00) - -
IDAS - (0.00) - (0.00) 92.35 (0.00) 81.36 (0.00) 72.31 (0.00) 37.30 (0.00) 75.74 (0.00) 63.01 (0.00) 25.57 (0.00) 30.61 (0.00) - -
SPILL 85.01 (0.29) 71.05 (0.83) 93.77 (0.35) 85.14 (1.04) 72.65 (0.32) 37.11 (1.62) 77.62 (0.46) 62.42 (2.06) 32.78 (0.52) 33.21 (2.33) 72.37 57.79
TWIST (Ours) 85.46 (0.18) 74.66 (1.01) 93.92 (0.09) 85.62 (0.71) 72.34 (0.28) 38.40 (0.81) 76.43 (0.03) 63.17 (0.25) 30.62 (0.24) 37.04 (1.55) 71.75 59.78
Table 5: Five-benchmark results (Emerging Data Scenario, K-means). Averages over 5 runs (±SD). Gemma is used in TWIST. Plain: direct embedding clustering. Bold numbers: best within the embedder. SCCL/ClusterLLM/LLMEdgeRefine cited directly from prior work. SCCL results are cited from ClusterLLM; IDAS results are cited from LLMedgeRefine.
GCD scenario
Table 6 shows the HDBSCAN results for GCD. Our method consistently improves the performance of the existing embedder. Table 7 shows the K-means results for GCD. Our approach is on average the best in the unlabeled setting, without any known categories. It even outperforms baselines that know $10\%$ of the categories, and is comparable to other baselines that know $25\%$ of the categories (or up to $75\%$ in the case of IntentGPT). These strong results, together with the fewer assumptions required by our method, highlight its effectiveness and efficiency. Notably, BANK77 and CLINC150 perform similarly with both clustering algorithms, while Stackoveflow shows a noticeable drop.
Bank 77 Clinc150 StackO Average
KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
- Plain 78.78 60.39 83.02 51.82 66.26 57.20 76.02 56.47
$\mathbf{0\%}$ TWIST 87.43 75.16 95.19 89.07 78.54 71.70 87.05 78.64
Table 6: Three-benchmark results (GCD scenario, HDBSCAN) Gemma is used in TWIST. Bold numbers: best. KCR ratio of known category. All-MiniLM-L6-v2 used as backbone embedder. Plain: direct embedding clustering. StackO: StackOverflow dataset.
Bank 77 Clinc150 StackO Average
KCR Method NMI Acc NMI Acc NMI Acc NMI Acc
- Plain 79.08 61.27 89.46 73.73 72.49 72.68 80.34 69.08
$\mathbf{0\%}$
IDAS 82.84 67.43 93.82 85.48 81.26 83.82 85.97 78.91
IntentGPT 81.42 64.22 94.35 83.20 - - - -
TWIST (Ours) 87.67 77.10 95.21 89.07 81.69 82.56 88.19 82.91
$\mathbf{10\%}$
Loop 79.14 64.97 93.52 84.89 75.98 80.50 82.88 76.79
GLEAN 82.23 67.99 95.21 88.71 79.67 82.40 85.70 79.70
$\mathbf{25\%}$
ALUP 84.06 74.61 94.84 88.40 76.58 82.20 85.16 81.74
Loop 83.37 71.40 94.38 86.58 79.10 82.20 85.62 80.06
GLEAN 85.62 76.98 96.27 91.51 80.90 84.10 87.60 84.20
$\mathbf{75\%}$ IntentGPT 85.94 77.21 96.06 88.76 - - - -
Table 7: Three-benchmark results (GCD scenario, K-means). Averages over 5 runs with Gemma. KCR: known categories ratio. All-MiniLM-L6-v2 used as backbone embedder in TWIST. Loop results are cited from GLEAN. Plain: direct embedding clustering.
6.2 Further analysis
Effect of Different LLMs
Our main results were obtained with gemma-2-9b-it. To investigate whether using a different LLM affects performance, Figure 2 shows the results obtained with all-MiniLM-L6-v2 as the embedder and three different LLMs (see Appendix E for details). Overall, the differences between Qwen3-8B and gemma-2-9b-it are similar but an obvious drop with Llama3.1-8B-Instruct. Because Qwen3-8B achieves comparable results, we use it along with all-MiniLM-L6-v2 for remaining analyses to reduce computational cost.
Figure 2: ACC and NMI across all test datasets (Emerging Data and GCD) for three LLMs. Results for K-means are averaged over 5 runs, HDBSCAN from a single run.
Effect of the Number of Closest Neighbors m and Iterative Updates
We use $m=30$ across all the experiment, but this raises the question of whether different values of $m$ would affect performance. Figure 3(a) shows that using too few neighbors leads to lower results. To investigate the effect of iterative updates, Figure 3(b) shows that in subsequent iterations, the results substantially improve, especially after the first iteration. Since most sparse vectors converge within 5 iterations, there are only minimal changes afterward: We find that most datasets have 90% convergence ratio by the 5th iteration (see Appendix F).
Figure 3: Comparison of (a) varying negihbor $m$ and (b) iterative updates on ACC and NMI across all test datasets (Emerging Data and GCD).
A scalable implementation of our method
Our representation relies on LLM guidance for each text, which can be computationally expensive when the dataset is extremely large. To assess scalability, we additionally explore clustering over the full dataset. The goal is to implement our approach on a subset of the texts and then use a model to generalize to the remaining texts (a form of knowledge distillation), without requiring further LLM computation. Specifically, we train a lightweight MLP ($\sim$10M parameters) to map pre-trained embeddings to sparse vectors. See Appendix G for experiment details. Note that we do not use any labeled data. Table 8 shows that our approach still improves the original embeddings by a large margin, which proves the scalability of our method, offering substantial savings in time and computational resources.
7 Conclusions
We propose TWIST (Short Text Clustering through Iterative Sparse Vector Updating with LLMs), an intuitive, domain-adaptive, label-free, and training-free method for short text clustering. Compared with other SOTA methods, TWIST is built under minimal assumptions and requirements. TWIST achieves better or similar results to state-of-the-art approaches without assuming prior knowledge of clusters or labels, making it more aligned with real-world scenarios. Experiments on diverse datasets and smaller LLMs show that our method is robust in low-resource settings, and also scalable to large datasets. In the future, we plan to expand our approach to clustering in specialized domains and explore ways to incorporate human feedback into our method pipeline.
Bank 77 Clinc150 StackOverf Average
Method NMI Acc NMI Acc NMI Acc NMI Acc
Kmeans
Plain 76.31 59.41 88.99 73.99 68.42 65.93 77.91 66.44
TWIST (Ours) 83.80 73.58 92.78 87.27 72.98 76.90 83.19 79.25
HDBSCAN
Plain 76.15 60.27 84.31 62.84 63.01 61.80 74.49 61.64
TWIST (Ours) 83.39 71.79 92.55 86.04 71.08 69.58 82.34 75.80
Table 8: Clustering results on the entire dataset using knowledge distillation. Averages over 5 runs for K-means. Plain: Directly embedding the test dataset with All-MiniLM-L6-v2.
Acknowledgments
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
Limitations
Language. Like most of the prior work, we only focus on English utterance datasets. This is relevant because most of benchmark datasets are in English. Our method is completely data-driven and not dependent on labeled data and it is therefore applicable to other languages than English.
LLM capability. Our method relies on the abilities of LLMs. Although we empirically show that small LLMs can achieve good results, this is not guaranteed when the dataset comes from a rare domain.
Embedder capability. Although we empirically show that our approach generally improves performance compared to directly using the embedder for clustering, if the embeddings provide a poor initial result, the improvement might be limited.
Appendix A Comparison of Data Splits in GCD scenario: Our Method vs. Prior Work
Table 9 shows the train–validation–test splits used in Loop (An et al., 2024), ALUP (Liang et al., 2024b) and GLEAN Zou et al. (2025b) for the three benchmark datasets. In their setting, a specified ratio of all categories is randomly selected as known categories, referred to as the known category ratio (KCR). For each known category, 10% of the data is chosen to form the labeled dataset [^2] IntentGPT (Rodriguez et al., 2024) mention they use fewer than 10% for each known category, but they do not specify the exact percentage., while the remaining samples constitute the unlabeled dataset. Together, these two subsets form the training dataset. In our experiments, we do not use the validation set and use only 30% of the training dataset, restricted to the unlabeled setting with no access to labeled data. Evaluation is conducted on the same test split.
They train an embedder using the training split and optimizes it based on the labeled validation set, with the test set used solely for evaluation. In contrast, our method performs test-time embedding construction, so it requires to access test data, similar to IDAS De Raedt et al. (2023) and IntentGPT Rodriguez et al. (2024), using all available text data ($D=D^{test}\cup D^{u}$) to construct embeddings.
Since our clustering is unsupervised, no label information is used. In practice, clustering methods require access to the text to be clustered in order to construct representations, and our approach naturally uses all available texts.
# clusters # train # valid # test # total
Bank77 77 9,003 1,000 3,080 13,083
Clinc150 150 18,000 2,250 2,250 22,500
StackO 20 18,000 1,000 1,000 20,500
Table 9: Train–Validation–Test splits from previous studies (An et al., 2024; Liang et al., 2024b; Zou et al., 2025b). ‘StackO’ is the StackOverflow dataset
Appendix B Selection of $d$
Table 10 shows the two datasets that we used for selecting the dimensionality of the sparse vector $d$. Figure 4 shows that $1024$, $2048$, and $\min\{4096,N\}$ give similar performance for both K-means and HDBSCAN. We selected value $d=2048$ and used this value across all datasets and experiments.
[…]
ArxivS2S 93 3,674
Reddit 50 3,217
Table 10: External datasets used for selecting the dimensionality hyperparameter $d$
Figure 4: The average ACC and NMI across two datasets for different $d$. Results for K-means are averaged over 5 runs, while HDBSCAN is shown from a single run. $4096$ here means $\min\{4096,N\}$ because the dataset size can be smaller.
Appendix C Prompts for Instructor embedder
Our method is built on top of existing embedders. For instructor-large, a prompt can be added to generate task-specific embeddings. We follow previous research (Su et al., 2022; Liang and Liao, 2023; Lin et al., 2025) and use mostly the same prompts. The prompts we used for the different datasets are listed in Table 11.
Instructor:
Bank77: Represent the bank purpose for retrieval:
Mtop and Stackoverflow: Represent the sentence for retrieval:
Clinc150 and Massive: Represent the sentence for retrieving the purpose:
GoEmo: Represent an emotion sentence for retrieval:
Table 11: The prompts used in Instructor.
Appendix D Estimated Number of Clusters $\hat{K}$ by HDBSCAN
Table 12 and Table 13 show the estimated number of clusters, $\hat{K}$, for the two scenarios. The results indicate that the refined embeddings produced by our method are generally closer to the ground truth than the pre-trained (Plain) embeddings. An exception is the Bank77 dataset.
Bank 77 Clinc150 Mtop Massive GoEmo
Bert
Plain 2 111 62 2 2
TWIST 73 154 80 57 63
MiniLM
Plain 78 139 17 27 11
TWIST 68 150 62 56 44
E5
Plain 73 134 9 2 3
TWIST 70 142 82 57 58
Instructor
Plain 71 144 18 29 2
TWIST 84 148 99 58 57
True $K$ 77 150 102 59 27
Table 12: Emerging scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. gemma-2-9b-it used in TWIST.
Bank 77 Clinc150 Stackoverflow
Plain 78 82 13
TWIST 84 142 30
True $K$ 77 150 20
Table 13: GCD scenario: Estimated Number of Clusters $\hat{K}$ by HDBSCAN. All-MiniLM-L6-v2 used as backbone embedder. gemma-2-9b-it was used in TWIST.
Bank 77 Clinc150 Mtop Massive GoEmo Bank 77GCD Clinc150GCD StackOGCD Average
NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc NMI Acc
Kmeans
Gemma 85.57 74.07 93.65 86.35 71.92 37.83 76.18 66.15 20.69 24.44 87.67 77.10 95.21 89.07 81.69 82.56 76.57 67.20
Qwen 85.13 73.18 93.59 86.96 72.23 39.03 74.55 63.64 20.12 22.80 87.04 75.78 94.80 88.21 80.69 82.32 76.02 66.49
Llama 82.46 69.81 91.03 82.21 70.08 36.21 72.47 60.47 17.60 19.67 84.34 70.91 93.83 86.27 75.69 78.68 73.44 63.03
HCBSCAN
Gemma 85.61 75.29 93.86 86.44 74.52 50.36 76.67 67.51 21.63 23.27 87.43 75.16 95.19 89.07 78.54 71.70 76.68 67.35
Qwen 85.04 73.90 93.56 86.42 74.42 50.11 74.39 62.36 22.10 19.81 86.81 75.29 94.78 88.18 78.07 71.50 76.15 65.95
Llama 82.45 68.70 91.00 78.64 72.80 51.12 73.19 63.84 18.53 16.84 84.16 69.35 93.84 84.18 74.78 72.20 73.84 63.11
Table 14: Further analysis of results using different LLMs.
Appendix E Results by Different LLMs
We report the detailed results of all datasets by different LLMs in Table 14. Qwen and Gemma consistently outperform LLaMA. In addition, the average differences between K-means and HDBSCAN are minimal across both metrics for each LLM.
Appendix F Percentage of texts converged at iteration $t$
Figure 5 shows ratio of text convergence. It shows most datasets have 90% convergence ratio by the 5th iteration, and by the 7th iteration, all datasets achieve a 95% convergence ratio.
Figure 5: Ratio of texts converged at iteration $t$. All-MiniLM-L6-v2 (as embedder) and Qwen3-8B (as LLM) were used.
Appendix G Details of the full dataset clustering
Datasets Used with LLM Guidance
Table 15 shows the number of examples we use with the LLM. The number of examples with LLM guidance is the same as in Table 1, i.e. $|D^{test}|+|D^{u}|$, so we do not need to re-run all experiments. The examples account for 29.6% to 44.2% of datasets. Examples with LLM guidance are used to train the MLP, since their sparse vectors have been derived. The pre-trained embeddings from all-MiniLM-L6-v2 serve as input, and the LLM-generated sparse vectors serve as target output for supervised learning with an MLP. 20% of the examples are used as a validation set to tune the hyperparameters.
MLP Architecture and Hyperparameter Search
We build an MLP for each dataset. We only tune the hidden layer dimensionality using the validation set, with search over $\{512,1024,2048\}$.
Clustering of the entire dataset with MLP output
We first encode all examples into pre-trained embeddings using all-MiniLM-L6-v2, then use the trained MLP to derive sparse versions of these embeddings. These sparse vectors are used to cluster all examples.
Dataset $K$ Total LLM No-LLM LLM ratio
Bank77 77 13,083 5,780 7,303 44.2%
Clinc150 150 22,500 7,650 14,850 29.6%
Stackoverflow 20 20,000 6,400 13,600 32.0%
Table 15: Dataset Statistics.
Evaluation
For HDBSCAN, we set min_samples = 1 to reduce noise. The noise points are assigned to the nearest cluster centroid. We perform an incremental grid search over min_cluster_size values of $50,75,100,125,\ldots,250$ – five times the original range – stopping early if the silhouette score does not improve. For K-means, we set the number of clusters equal to the ground-truth cluster count $K$.